Car A, moving at 16.64 m/s, is 1,374.39 m ahead of car B which is moving at 34.46 m/s. Car B decides to pass car A and accelerates at 2.41 m/s2. How long will it take car B to overtake car A?
Answer #1

velocity of Car A= 16.64 m/s

velocity of Car B= 34.46 m/s

the distance between Cars is 1.374 km,

hence relative to Car A, the Car B moves 34.46-16.64= 17.82 m/s more

then Car B accelerates 2.41 m/s2,

so the final velocity of the Car B is, v= v0+at

where t=1, v0= 34.46 m/s, then v= 34.46+2.41= 36.87 m/s in 1 sec time.,

and the Car A is moving with 16.64 m/s so the resultant velocity is, 36.87-16.64= 20.23

so to cross 1.3 km, time= 1374/20.23= 67 seconds it need